# Workflow as a Service Broker in Cloud Environment: A Systematic Literature Review

### v1

@@ 4.2.7. Scheduling Algorithm
Multi-criteria decision-making (MCDM) is a field that focuses on choosing the best alternative from several options by evaluating multiple criteria or attributes (Taherdoost and Madanchian, 2023). Among the popular MCDM methods, only the Technique for Order of Preference by Similarity to Ideal Solution (TOPSIS) has been applied to the MWSP (Xia et al., 2023; Chakravarthi and Shyamala, 2021; Chakravarthi et al., 2022). Chakravarthi et al. (Chakravarthi and Shyamala, 2021; Chakravarthi et al., 2022) use TOPSIS to determine the best VM for a ready task in a dynamic scheduling algorithm. They input a $m\times r$ decision matrix into TOPSIS, where $m$ represents the number of available VMs, and $r$ is the number of criteria (3 in this case): execution time, transfer time, and execution cost. Another approach is Game theory, which is used to model scenarios where multiple independent players interact and influence each other’s outcomes (Nisan et al., 2007). Among the reviewed studies, only one uses this method. Wu and Yuandou (Wu and Yuandou, 2018) apply multi-stage dynamic game theory to minimize makespan while enhancing fairness and utilization. Each optimization objective—makespan, fairness, and utilization—is considered as a competing player. The Nash equilibrium is determined as the optimal solution to the problem.
@@
Finally, mathematical optimization (or mathematical programming) utilizes techniques like Linear/Nonlinear Programming (LP/NLP) to address optimization problems. Karmakar et al. (Karmakar et al., 2024) apply NLP to minimize execution costs for a set of deadline-constrained workflows. They first identify the critical paths within the workflow and then aim to minimize the total required MIPS for tasks on these critical paths, ensuring the workflow’s deadline is met. The problem is formulated as an NLP, and an approximate solution is found using the Lagrange Multiplier Method. Figure 6(b) shows the distribution of articles for each algorithm. It is clear that heuristic algorithms are the most frequently used, followed by meta-heuristic algorithms.
4.3. Resource provisioner taxonomy
@@
[…]
Figure 7. Resource provisioner taxonomy.
@@
4.3.1. Resource Type
As shown in Figure 8(a), the majority of the articles reviewed schedule workflows directly on VMs, which are traditional IaaS resources (Chen et al., 2020b). However, containers have recently emerged as a lightweight virtualization technology, allowing users to bundle tasks as standalone, isolated applications. Consequently, some articles have adopted containers for task scheduling on VMs. For example, in (Taghavi et al., 2023; Li et al., 2022c), the broker rents large VMs with numerous CPU cores and ample memory, using containers to bundle and execute multiple tasks on a single VM. Although serverless computing and FaaS have emerged as promising models for executing scientific workflows in the cloud (Malawski et al., 2020), only Sun et al. (Sun et al., 2024) have explored the scheduling of multiple workflows using a combination of serverless functions and VMs. This combination shows promise and will be discussed further in future directions (Section 5).
@@ 4.3.1. Resource Type
Figure 8. Number of articles for different resource types and pricing
4.3.2. Resource Diversity
@@
Most of the papers reviewed focused on a cloud provider offering heterogeneous resources with varying configurations (such as CPU cores, memory, and storage), aiming to identify the optimal resource configuration for each task based on its requirements and scheduling objectives and constraints (e.g., deadlines and budgets). However, some researchers opted for a homogeneous environment, using a single resource type for all tasks, typically to simplify the scheduling process and reduce the search space (Liu et al., 2024; Zhu et al., 2020). To justify this approach, Malawski et al. (Malawski et al., 2015) argue that, for most applications, there will generally be one or two VM types that offer the best price/performance ratio. In contrast, Taheri et al. (Taheri et al., 2023) utilized only a single type of large VMs, then employed containers to schedule tasks on these VMs.
4.3.3. Resource Pricing
As depicted in Figure 8(b), the majority of the reviewed articles utilized the widely known on-demand pricing model for executing users’ workflows, leveraging the elasticity of cloud resources (Adhikari and Koley, 2018; Shen et al., 2025). However, some studies, either implicitly or explicitly, considered a fixed pool of reserved resources, which doesn’t fully capture the essence of cloud computing (Pan and Wei, 2024; Zhang et al., 2023). Despite this, Chen et al. (Chen et al., 2020a) used reserved resources for executing periodic workflows, proposing a heuristic algorithm to determine the optimal amount of reserved resources. More interestingly, several studies explored how the WaaS broker can reduce execution costs by combining spot and on-demand instances (Taghavi et al., 2023; Zolfaghari and Abrishami, 2022; Sampaio and Barbosa, 2020; Chen et al., 2024b; Zhou et al., 2016), or reserved and on-demand instances (Taheri et al., 2023; Zhu et al., 2020).
@@ 4.3.3. Resource Pricing
Zolfaghari et al. (Zolfaghari and Abrishami, 2022) developed a multi-class workflow scheduling algorithm, where tasks are assigned to different resource classes based on their priority. Higher-priority tasks are sent to higher-class resources, which are reliable but expensive on-demand instances, while lower-priority tasks are allocated to cheaper spot instances with lower bid prices. If a task fails to execute within a specified time, it is moved to a higher-class resource until completion. Similarly, Zhou et al. (Zhou et al., 2016) combined spot and on-demand resources for workflow execution. They initially used the A* algorithm to identify the optimal on-demand instance type for each task to minimize costs while meeting the workflow’s deadline. Then, a refinement algorithm replaces some on-demand instances with spot instances to further reduce costs. Taheri et al. (Taheri et al., 2023) focused on scheduling periodic workflows using a mix of reserved and on-demand instances. Their scheduling algorithm prioritizes reserved instances for each task until its sub-deadline is reached. If resources are insufficient, the algorithm switches to on-demand instances to avoid missing the sub-deadline. Combining different pricing models to reduce execution costs is a promising strategy that will be discussed in future directions (Section 5).
4.3.4. Provision Timing
@@
Provision timing is categorized into as-needed and periodic approaches. In as-needed provisioning, the broker acquires new resources when a task (or a group of tasks) requires them—typically because they cannot be scheduled on the currently available resources without violating an optimization criterion or constraint (e.g., a sub-deadline). In this approach, the resource provisioner is usually integrated within the task scheduler rather than being a separate module. As a result, the scheduling algorithm is responsible for both provisioning and de-provisioning resources. Whenever there is a resource shortage preventing the scheduling of a task, the algorithm triggers commands to lease new resources. The key advantage of this method is its prompt response to environmental events. For instance, in reference (Taheri et al., 2023), the scheduling algorithm calculates a parameter called laxity for each task, indicating the amount of time it can be delayed before missing its sub-deadline. If a task is ready to execute but no idle resources are available, the scheduler postpones it until the necessary resources become free, provided its laxity is still positive. If the required resources are not released and the laxity reaches zero, the scheduler proceeds to request new resources.
@@
In contrast, periodic provisioning, commonly used in dynamic scheduling algorithms (whether immediate or periodic), requires the task scheduler to allocate workflow tasks to currently available resources. The resource provisioner operates as an independent module, periodically assessing the current resources based on parameters such as CPU load, task queue size, and other metrics. If necessary, the provisioner decides whether to lease additional resources or release some existing ones. One disadvantage of this method, compared to the as-needed approach, is that it responds to changes in the environment and workflow requirements later, typically at the end of each time interval. However, this delay can also be an advantage. In the as-needed approach, whenever a task requests a resource, it is immediately granted, often without considering the status of other tasks (either upcoming or current). This can result in prematurely leasing resources when, in a few moments, a suitable resource might become available, or it might be better to wait and request more powerful resources as new tasks arrive. By delaying resource provision until the end of each time interval, a more informed decision can be made, taking into account the overall status of resources and the tasks in the queue. However, this approach risks missing certain constraints, such as deadlines, so the broker must make more conservative decisions in these situations.
@@
Taghavi et al. (Taghavi et al., 2023) combine dynamic periodic scheduling with periodic provisioning, where the provisioning intervals are significantly longer than the scheduling intervals (ranging from 5 to 10 times longer). During each provisioning interval, the provisioner checks whether the number of idle cores on the VMs is sufficient to accommodate the required cores for the ready tasks. If not, new VMs are leased from the cloud provider. In contrast, Malawski et al. (Malawski et al., 2015) implement dynamic immediate scheduling with periodic provisioning. Their approach computes resource utilization at the end of each interval and adjusts the number of leased VMs if the utilization exceeds or falls below certain thresholds. Tarafdar et al. (Tarafdar et al., 2023) employ a different approach to periodic provisioning, where, at the end of each time interval, they attempt to predict the number of workflows arriving in the next interval using historical data from previous intervals and scale the resources up or down accordingly.
4.3.5. Provisioning Delay
@@
When a broker requests the provisioning of a VM from a cloud provider, it does not become available immediately. Instead, it takes a significant amount of time to be deployed on a physical host and for its operating system to boot. This period is referred to as provisioning delay. Even though containers, a lightweight virtualization technology, have a much shorter provisioning delay than VMs, they still experience some delay. This provisioning delay can be substantial, even in the case of containers, and thus, to make accurate provisioning and scheduling decisions, this parameter should be considered. In particular, provisioning delay plays a crucial role in provisioning algorithms, as in some situations it may be better not to release an idle resource so that it can be used for future tasks, avoiding the delay associated with provisioning new resources. To minimize VM provisioning delays, some researchers opt to use larger VMs and run multiple tasks on these VMs using container technology (Taghavi et al., 2023). However, only about 32% of the reviewed articles considered provisioning delay (Partheeban and Kavitha, 2019; Li et al., 2022d), which we refer to as delay-aware, while the remainder of the articles did not mention it at all (Ye et al., 2023; Li et al., 2024a), and are categorized as instantaneous.
4.3.6. Deployment Model
@@
This taxonomy examines how the broker acquires its required resources, either from a single source or multiple sources. The majority of articles (66 papers) utilize a single-cloud provider to lease their resources. However, relying on a single provider has its drawbacks, including vendor lock-in, limited resource diversity, restricted pricing models, provider unavailability, and geographical location constraints (Khorasani et al., 2024). As a result, some studies (7 papers) adopt the multi-cloud model, where the broker leases resources from multiple, independent cloud providers (Taghinezhad-Niar and Taheri, 2023; Xu et al., 2023; Chen et al., 2020b; Sun et al., 2024). In contrast, a few articles (3 papers) implement a hybrid-cloud approach, where the broker uses resources from both a local private cloud and a public cloud, turning to the latter when local resources are insufficient (Li et al., 2022d; Sun et al., 2023; Sharif et al., 2014). References (Sun et al., 2023; Sharif et al., 2014) specifically use the private cloud to meet privacy requirements, where certain workflow tasks must be executed on a private cloud due to the sensitive nature of the data involved.
4.4. Monitoring and fault tolerance taxonomy
@@
[…]
Figure 9. Monitoring and fault tolerance taxonomy.
@@
Resource failures are an inevitable aspect of computer datacenters, making it crucial for WaaS brokers to monitor resource performance and respond appropriately when failures occur. However, this requirement has not been adequately addressed in many reviewed studies. This is especially important in the spot pricing model, where the unreliability of resources and the potential for service providers to reclaim them during task execution are concerns. Despite these challenges, only 11 articles, including all five that utilize spot instances, address fault tolerance.
@@
Figure 9 outlines a taxonomy for monitoring and fault tolerance. The simplest method, re-execution, involves transferring tasks from a failed resource to a new one, but this results in the loss of all progress made up to that point (Malawski et al., 2015; Esteves and Veiga, 2015; Sampaio and Barbosa, 2020; Zhou et al., 2016). This can be costly for the WaaS broker, particularly for scientific workflows with long-running tasks and strict deadlines. To minimize this cost, some papers use checkpointing (Taghavi et al., 2023; Zhou et al., 2024; Zolfaghari and Abrishami, 2022; Xu et al., 2023), where snapshots of tasks or VMs are periodically saved. If a failure occurs, the task is reassigned to a new instance, and the most recent snapshot is restored. Additionally, some cloud providers like AWS offer a two-minute notice before stopping or terminating spot instances, which allows enough time to take a snapshot. However, the process of taking and restoring snapshots can be time-consuming.

### v2

@@ 4.2.7 Scheduling Algorithm
Multi-criteria decision-making (MCDM) is a field that focuses on choosing the best alternative from several options by evaluating multiple criteria or attributes [162]. Among the popular MCDM methods, only the Technique for Order of Preference by Similarity to Ideal Solution (TOPSIS) has been applied to the MWSP [185, 29, 28]. Chakravarthi et al. [29, 28] use TOPSIS to determine the best VM for a ready task in a dynamic scheduling algorithm. They input a $m\times r$ decision matrix into TOPSIS, where $m$ represents the number of available VMs, and $r$ is the number of criteria (3 in this case): execution time, transfer time, and execution cost.
Game theory is another approach used to model scenarios in which multiple independent players interact and influence each other’s outcomes [123]. Among the reviewed studies, only one uses this method. Wu and Yuandou [184] apply multi-stage dynamic game theory to minimize makespan while enhancing fairness and utilization. Each optimization objective—makespan, fairness, and utilization—is considered as a competing player. The Nash equilibrium is determined as the optimal solution to the problem.
@@
Finally, mathematical optimization (or mathematical programming) utilizes techniques like Linear/Nonlinear Programming (LP/NLP) to address optimization problems. Karmakar et al. [87] apply NLP to minimize execution costs for a set of deadline-constrained workflows. They first identify the critical paths within the workflow and then aim to minimize the total required MIPS for tasks on these critical paths, ensuring the workflow’s deadline is met. The problem is formulated as an NLP, and an approximate solution is found using the Lagrange Multiplier Method. Figure 6(b) shows the distribution of articles for each algorithm. It is clear that heuristic algorithms are the most frequently used, followed by meta-heuristic algorithms.
4.3 Resource Provisioner Taxonomy
@@
Figure 7: Resource provisioner taxonomy.
@@
4.3.1 Resource Type
As illustrated in Figure 8(a), the majority of reviewed articles schedule workflow tasks directly on VMs [42], which represent traditional IaaS resources. However, VMs are heavyweight, and their high provisioning delay can significantly reduce the efficiency of scheduling algorithms—especially in deadline-constrained scenarios, where such delays may result in missed deadlines. To mitigate provisioning overhead, brokers often keep idle VMs available for future use, which leads to reduced resource utilization—particularly in the case of short- and medium-duration tasks. In response to these limitations, containers have recently gained popularity as a lightweight virtualization technology. Containers offer near-instant provisioning and allow users to encapsulate tasks as isolated, self-contained applications. As a result, some studies have adopted containers to schedule multiple tasks on the same VM, improving utilization while avoiding provisioning delays. For example, in [159, 102], the broker rents large VMs with many CPU cores and substantial memory, then uses containers to bundle and execute several tasks concurrently on a single VM.
@@ 4.3.1 Resource Type
Figure 8: Number of articles for different resource types and pricing
Although serverless computing and FaaS have emerged as promising models for executing scientific workflows in the cloud [113], only Sun et al. [157] have investigated the scheduling of multiple workflows using a hybrid approach that combines serverless functions and VMs. Their study was conducted in a multi-cloud environment spanning AWS, GCP, and Azure. They applied GPHH hyper-heuristic (see Section 4.2.7) to automatically generate three heuristic rules for: (1) task selection, (2) CSP selection, and (3) resource selection. The third rule accounts for both VMs and serverless functions when selecting resources to execute tasks, aiming to minimize the total execution cost while satisfying deadline constraints. The combination of serverful resources and serverless functions shows significant potential, which is discussed as a promising future direction in Section 5.5.
To date, no prior work has addressed the scheduling of multiple workflows in hybrid quantum/classical environments, which is unsurprising given the early developmental stage of quantum computing and quantum algorithms. The sole exception is Qonductor [60], a WMS designed specifically for multiple hybrid workflows. However, as of mid-2025, it remains an arXiv preprint and has not appeared in a peer-reviewed journal. We discuss this emerging and promising area of hybrid scientific workflows, with particular attention to Qonductor, in the future directions outlined in Section 5.6.
4.3.2 Resource Diversity
@@
The majority of reviewed studies considered CSPs offering heterogeneous resources with varying configurations, such as CPU cores, memory, and storage, aiming to select the optimal resource configuration for each task based on its requirements and scheduling objectives. Conversely, some researchers have opted for a homogeneous environment, using a single resource type for all tasks to simplify scheduling and reduce the search space [111, 201]. Malawski et al. [114] justify this approach by noting that, for most applications, typically one or two VM types provide the best price/performance ratio. Similarly, Taheri et al. [163] utilized a single type of large VM with sufficient resources and employed containers to schedule multiple tasks on the same VM.
4.3.3 Resource Pricing
As shown in Figure 8(b), most of the reviewed studies employed the widely used on-demand pricing model to execute users’ workflows, taking advantage of the elasticity offered by cloud resources [4, 151]. In contrast, some studies, either implicitly or explicitly, relied on a fixed pool of reserved resources, which does not fully reflect the dynamic nature of cloud computing [125, 191]. Nevertheless, Chen et al. [37] utilized reserved resources for executing periodic workflows and proposed a heuristic algorithm to determine the optimal number of reserved instances, leveraging the flexible nature of reserved resources.
@@ 4.3.3 Resource Pricing
More interestingly, several studies explored how the WaaS broker can reduce execution costs by combining spot and on-demand instances [159, 202, 141, 38, 197, 109], or reserved and on-demand instances [163, 201]. Zolfaghari et al. [202] developed a multi-class workflow scheduling algorithm, where tasks are assigned to different resource classes based on their priority. Higher-priority tasks are sent to higher-class resources, which are reliable but expensive on-demand instances, while lower-priority tasks are allocated to cheaper spot instances with lower bid prices. If a task fails to execute within a specified time, it is moved to a higher-class resource until completion. Similarly, Zhou et al. [197] combined spot and on-demand resources for workflow execution. They initially used the A* algorithm to identify the optimal on-demand instance type for each task to minimize costs while meeting the workflow’s deadline. Then, a refinement algorithm replaces some on-demand instances with spot instances to further reduce costs. Taheri et al. [163] focused on scheduling periodic workflows using a mix of reserved and on-demand instances. Their scheduling algorithm prioritizes reserved instances for each task until its sub-deadline is reached. If resources are insufficient, the algorithm switches to on-demand instances to avoid missing the sub-deadline. The combination of different pricing models to reduce execution costs represents a promising strategy, which is further discussed in the future directions in Section 5.3.
4.3.4 Provision Timing
@@
Provision timing is categorized into as-needed and periodic approaches. In as-needed provisioning, the broker acquires new resources when a task (or a group of tasks) requires them—typically because they cannot be scheduled on the currently available resources without violating an optimization criterion or constraint (e.g., a sub-deadline). In this approach, the resource provisioner is usually integrated within the task scheduler rather than being a separate module. Whenever there is a resource shortage preventing the scheduling of a task, the algorithm triggers commands to lease new resources. The key advantage of this method is its prompt response to environmental events. For instance, in reference [163], the scheduling algorithm calculates a parameter called laxity for each task, indicating the amount of time it can be delayed before missing its sub-deadline. If a task is ready to execute but no idle resources are available, the scheduler postpones it until the necessary resources become free, provided its laxity is still positive. If the required resources are not released and the laxity reaches zero, the scheduler proceeds to request new resources.
@@
In contrast, periodic provisioning, commonly used in dynamic scheduling algorithms, requires the task scheduler to allocate workflow tasks to currently available resources, without actively managing resource provisioning. The resource provisioner operates as an independent module, periodically assessing the current resources based on parameters such as CPU load, task queue size, and other metrics. If necessary, the provisioner decides whether to lease additional resources or release some existing ones. One drawback of this approach is that it reacts to changes in the environment and workflow requirements with a delay, typically only at the end of each time interval. This lag can risk violating certain constraints, such as deadlines, forcing the broker to adopt more conservative decisions in some cases. On the other hand, this delay can also be advantageous. In the as-needed approach, resources are granted immediately whenever a task requests them, often without considering the state of other running or pending tasks. This can lead to premature resource allocation, whereas shortly thereafter a suitable resource might become available, or it might be preferable to wait and rent more powerful resources as new tasks arrive. By deferring resource provisioning until the end of each interval, the broker can make more informed decisions, taking into account the overall status of both resources and queued tasks.
@@
Taghavi et al. [159] combine dynamic periodic scheduling with periodic provisioning, where the provisioning intervals are significantly longer than the scheduling intervals (ranging from 5 to 10 times longer). During each provisioning interval, the provisioner checks whether the number of idle cores on the VMs is sufficient to accommodate the required cores for the ready tasks. If not, new VMs are leased from the CSP. In contrast, Malawski et al. [114] implement dynamic immediate scheduling with periodic provisioning. Their approach computes resource utilization at the end of each interval and adjusts the number of leased VMs if the utilization exceeds or falls below certain thresholds. Tarafdar et al. [165] employ a different approach to periodic provisioning, where, at the end of each time interval, they attempt to predict the number of workflows arriving in the next interval using historical data from previous intervals and scale the resources up or down accordingly.
4.3.5 Provisioning Delay
@@
When a broker requests the provisioning of a VM from a CSP, it does not become available immediately. Instead, it takes a significant amount of time to be deployed on a physical host and for its operating system to boot. This period is referred to as provisioning delay. Even though containers, a lightweight virtualization technology, have a much shorter provisioning delay than VMs, they still experience some delay. This provisioning delay can be substantial, even in the case of containers, and thus, to make accurate provisioning and scheduling decisions, this parameter should be considered. In particular, provisioning delay plays a crucial role in provisioning algorithms, as in some situations it may be better not to release an idle resource so that it can be used for future tasks, avoiding the delay associated with provisioning new resources. To minimize VM provisioning delays, some researchers opt to use larger VMs and run multiple tasks on these VMs using container technology [159]. However, only about 30% of the reviewed articles considered provisioning delay [126, 104], which we refer to as delay-aware, while the remainder of the articles did not mention it at all [190, 99], and are categorized as instantaneous.
4.3.6 Deployment Model
@@
This taxonomy examines how the broker acquires its required resources, either from a single CSP or multiple CSPs. The majority of articles (76 papers) utilize a single-cloud provider to lease their resources. However, relying on a single CSP has its drawbacks, including vendor lock-in, limited resource diversity, restricted pricing models, provider unavailability, and geographical location constraints [90]. As a result, some studies (7 papers) adopt the multi-cloud model, where the broker leases resources from multiple, independent CSPs [161, 186, 42, 157, 196, 105, 184]. In contrast, a few articles (4 papers) implement a hybrid-cloud approach, where the broker uses resources from both a local private cloud and a public cloud, turning to the latter when local resources are insufficient [104, 156, 148, 67].
4.4 Monitoring and Fault Tolerance Taxonomy
@@
Figure 9: Monitoring and fault tolerance taxonomy.
@@
Resource failures are an inevitable aspect of computer datacenters, making it crucial for WaaS brokers to monitor resource performance and respond appropriately when failures occur. However, this requirement has not been adequately addressed in many reviewed studies. This is especially important in the spot pricing model, where the unreliability of resources and the potential for CSPs to reclaim them during task execution are concerns. Despite these challenges, only 13 articles, including all 6 that utilize spot instances, address fault tolerance.
@@
Figure 9 outlines a taxonomy for monitoring and fault tolerance. The simplest method, re-execution, involves transferring tasks from a failed resource to a new one, but this results in the loss of all progress made up to that point [114, 53, 141, 197]. This can be costly for the WaaS broker, particularly for scientific workflows with long-running tasks and strict deadlines. To minimize this cost, some papers use checkpointing [159, 198, 202, 186], where snapshots of tasks or VMs are periodically saved. If a failure occurs, the task is reassigned to a new instance, and the most recent snapshot is restored. Additionally, most CSPs provide advance notice before stopping or terminating spot instances, allowing sufficient time to take a snapshot. However, the process of taking and restoring snapshots can be time-consuming.